# Sparse Alignment Enhanced Latent Diffusion Transformer for Zero-Shot Speech Synthesis

### v1

@@ Abstract
While recent zero-shot text-to-speech (TTS) models have significantly improved speech quality and expressiveness, mainstream systems still suffer from issues related to speech-text alignment modeling: 1) models without explicit speech-text alignment modeling exhibit less robustness, especially for hard sentences in practical applications; 2) predefined alignment-based models suffer from naturalness constraints of forced alignments. This paper introduces S-DiT, a TTS system featuring an innovative sparse alignment algorithm that guides the latent diffusion transformer (DiT). Specifically, we provide sparse alignment boundaries to S-DiT to reduce the difficulty of alignment learning without limiting the search space, thereby achieving high naturalness. Moreover, we employ a multi-condition classifier-free guidance strategy for accent intensity adjustment and adopt the piecewise rectified flow technique to accelerate the generation process. Experiments demonstrate that S-DiT achieves state-of-the-art zero-shot TTS speech quality and supports highly flexible control over accent intensity. Notably, our system can generate high-quality one-minute speech with only 8 sampling steps. Audio samples are available at https://sditdemo.github.io/sditdemo/.
Sparse Alignment Enhanced Latent Diffusion Transformer
@@
Ziyue Jiang1, Yi Ren2, Ruiqi Li1, Shengpeng Ji1, Zhenhui Ye1, Chen Zhang2, Bai Jionghao1, Xiaoda Yang1, Jialong Zuo1, Yu Zhang1, Rui Liu3, Xiang Yin2, Zhou Zhao1 1Zhejiang University, 2ByteDance, 3Inner Mongolia University ziyuejiang341@gmail.com, zhaozhou@zju.edu.cn
@@ 1 Introduction
In recent years, neural codec language models (Wang et al., 2023; Zhang et al., 2023; Song et al., 2024; Xin et al., 2024) and large-scale diffusion models  (Shen et al., 2023; Matthew et al., 2023; Lee et al., 2024a; Eskimez et al., 2024; Ju et al., 2024; Yang et al., 2024d, b) have brought considerable advancements to the field of speech synthesis. Unlike traditional text-to-speech (TTS) systems (Shen et al., 2018; Jia et al., 2018; Li et al., 2019; Kim et al., 2020; Ren et al., 2019; Kim et al., 2021, 2022), these models are trained on large-scale, multi-domain speech corpora, which contributes to notable improvements in the naturalness and expressiveness of synthesized audio. Given only seconds of speech prompt, these models can synthesize identity-preserving speech in a zero-shot manner.
@@
  • Predefined alignment-based methods have prosodic naturalness constraints of forced alignments. During training, alignment paths (Ren et al., 2020; Kim et al., 2020) are directly introduced into their models (Matthew et al., 2023; Shen et al., 2023; Ju et al., 2024) to reduce the complexity of text-to-speech generation, which achieves higher intelligibility. Nevertheless, they suffer from the following two limitations: 1) predefined alignments constrain the model’s search space to produce more natural-sounding speech (Anastassiou et al., 2024; Chen et al., 2024a); 2) the overall naturalness is highly dependent on the performance of the duration model.
Intuitively, we can integrate the two aforementioned diffusion-based methods to pursue optimal performance. To be specific, we propose a novel sparse speech-text alignment strategy to enhance the latent diffusion transformer (DiT), termed S-DiT. In our approach, phoneme tokens are sparsely distributed within the corresponding forced alignment regions to provide coarse pronunciation information that is then refined by the latent DiT model. Experimental results demonstrate that S-DiT achieves nearly state-of-the-art speech intelligibility and speaker similarity on the LibriSpeech test-clean set (Panayotov et al., 2015) with only 8 sampling steps, while also exhibiting high speech naturalness. The main contributions of this work are summarized as follows:
  • We design a sparse alignment enhanced latent diffusion transformer (S-DiT) model, which effectively integrates the strengths of the two aforementioned speech-text alignment approaches. Notably, S-DiT also demonstrates greater robustness to duration prediction errors compared to methods with forced alignment.
@@
  • We successfully reduce S-DiT’s inference steps from 25 to 8 with the piecewise rectified flow (PeRFLow) technique, achieving highly efficient zero-shot TTS with minimal quality degradation. We also visualize the attention score matrices across various layers of S-DiT and obtain insightful findings in Appendix F.
@@ 2 Background
Figure 1: (a) The WaveVAE model; (b) Overview of S-DiT. We insert the sparse alignment anchors into the latent vector sequence to provide coarse alignment information. The transformer blocks in S-DiT will automatically build fine-grained alignment paths.
@@ 3 Method
This section introduces S-DiT. To begin with, we describe the architecture design of S-DiT. Then, we provide detailed explanations of the sparse alignment mechanism, the piecewise rectified flow acceleration technique, and the multi-condition classifier-free guidance strategy.
@@ 3.1 Architecture
We use the standard transformer block from LLAMA (Dubey et al., 2024) as the basic structure for S-DiT and adopt the Rotary Position Embedding (RoPE) (Su et al., 2024) as the positional embedding. During training, we randomly divide the latent vector sequence into a prompt region $z_{prompt}$ and a masked target region $z_{target}$, with the proportion of $z_{prompt}$ being $\gamma\sim U(0.1,0.9)$. We use $v_{\theta}$ to predict the masked target vector $\hat{z}_{target}$ conditioned on $z_{prompt}$ and the phoneme embedding $p$, denoted as $v_{\theta}(\hat{z}_{target}|z_{prompt},p)$. The loss is calculated using only the masked region $z_{target}$. S-DiT learns the average pronunciation from $p$ and the specific characteristics such as timbre, accent, and prosody of the corresponding speaker from $z_{prompt}$.
3.2 Sparse Alignment Enhanced Latent Diffusion Transformer (S-DiT)
In this subsection, we describe the sparse alignment strategy as the foundation of S-DiT, followed by the piecewise rectified flow and multi-condition CFG strategies to further enhance S-DiT’s capacity.
@@ 3.2 Sparse Alignment Enhanced Latent Diffusion Transformer (S-DiT)
Let’s first analyze the reasons behind the characteristics of different speech-text alignment modeling methods in depth. Implicitly modeling speech-text alignment is a relatively challenging task, which consequently leads to suboptimal speech intelligibility, particularly in hard cases. On the other hand, employing predefined hard alignment paths constrains the model’s search space to produce more natural-sounding speech. The characteristics of these systems motivate us to design an approach that combines the advantages of both: we first provide a rough alignment to S-DiT and then use attention mechanisms in Transformer blocks to construct the fine-grained implicit alignment path. The visualizations of the implicit alignment paths are included in Appendix F. In specific, denote the latent speech vector sequence as $z=[z_{1},z_{2},\cdots,z_{n}]$, the phoneme sequence as $p=[p_{1},p_{2},\cdots,p_{m}]$, and the phoneme duration sequence as $d=[d_{1},d_{2},\cdots,d_{m}]$, where $n$, $m$ is the length of the sequence. The length of the speech vector that corresponds to a phoneme $p_{i}$ is the duration $d_{i}$. Given $d=[2,2,3]$, the hard speech-text alignment path can be denoted as $a=[p_{1},p_{1},p_{2},p_{2},p_{3},p_{3},p_{3}]$. To construct the rough alignment $\tilde{a}$, we randomly retain only one anchor for each phoneme: $\tilde{a}=[\underline{M},p_{1},p_{2},\underline{M},\underline{M},\underline{M},P_{3}]$, where $\underline{M}$ represents the mask token. $\tilde{a}$ is downsampled with convolution layers to match the length of the latent sequence $z$. Then, we directly concatenate the downsampled $\tilde{a}$ and $z$ along the channel dimension. The anchors in $\tilde{a}$ provide S-DiT with approximate positional information for each phoneme, simplifying the learning process of speech-text alignment. At the same time, the rough alignment information does not limit S-DiT’s search space and also enables fine-grained control over each phoneme’s duration.
@@
We adopt Piecewise Rectified Flow (PeRFlow) (Yan et al., 2024) to distill the pretrained S-DiT model into a more efficient generator. Although our S-DiT is non-autoregressive in terms of the time dimension, it requires multiple iterations to solve the Flow ODE. The number of iterations (i.e., number of function evaluations, NFE) has a great impact on inference efficiency, especially when the model scales up further. Therefore, we adopt the PeRFlow technique to further reduce NFE by segmenting the flow trajectories into multiple time windows. Applying reflow operations within these shortened time intervals, PeRFlow eliminates the need to simulate the full ODE trajectory for training data preparation, allowing it to be trained in real-time alongside large-scale real data during the training process. Given number of windows $K$, we divide the time $t\in[0,1]$ into $K$ time windows $\{(t_{k-1},t_{k}]\}^{K}_{k=1}$. Then, we randomly sample $k\in\{1,\cdots,K\}$ uniformly. We use the startpoint of the sampled time window $z_{t_{k-1}}=\sqrt{1-\sigma^{2}(t_{k-1})}z_{1}+\sigma(t_{k-1})\epsilon$ to solve the endpoint of the time window $\hat{z}_{t_{k}}=\phi_{\theta}(z_{t_{k-1}},t_{k-1},t_{k})$, where $\epsilon\sim\mathcal{N}(0,I)$ is the random noise, $\sigma(t)$ is the noise schedule, and $\phi_{\theta}$ is the ODE solver of the teacher model. Since $z_{t_{k-1}}$ and $\hat{z}_{t_{k}}$ is available, the student model $\hat{\theta}$ can be trained via the following objectives:
@@
where $v_{\hat{\theta}}$ is the estimated drift force with parameter $\hat{\theta}$ and $t$ is uniformly sampled from $(t_{k-1},t_{k}]$. We provide details of PeRFlow training for S-DiT in Appendix C.
@@
where $c$ denotes the conditional state, $\varnothing$ denotes the unconditional state, and $\alpha$ is the guidance scale selected based on experimental results. Unlike standard classifier-free guidance, S-DiT’s conditional states $c$ consist of two components: phoneme embeddings $p$ and the speaker prompt $z_{prompt}$. In the experiments, as the text guidance scale increases, we observe that the pronunciation changes according to the following pattern: 1) starting with improper pronunciation; 2) then shifting to pronouncing with the current speaker’s accent; 3) and finally approaching the standard pronunciation of the target language. The detailed experimental setup is described in Appendix K. This allows us to use the text guidance scale $\alpha_{txt}$ to control the accent intensity. At the same time, the speaker guidance scale $\alpha_{spk}$ should be a relatively high value to ensure a high speaker similarity. Therefore, we adopt the multi-condition classifier-free guidance technique to separately control $\alpha_{txt}$ and $\alpha_{spk}$:
@@
S-DiT 0.3B LibriLight 0.71 0.78 1.82% 0.00 3.98 0.188
S-DiT-accelerated 0.3B LibriLight 0.70 0.78 1.86% -0.03 3.96 0.124
@@ 4.1 Experimental setup
We train S-DiT on the LibriLight (Kahn et al., 2020) dataset, which contains 60k hours of unlabeled speech derived from LibriVox audiobooks. All speech data are sampled at 16KHz. We transcribe the speeches using an internal ASR system and extract the predefined speech-text alignment using the external alignment tool (McAuliffe et al., 2017). We utilize three benchmark datasets: 1) the librispeech (Panayotov et al., 2015) test-clean set following NaturalSpeech 3 (Ju et al., 2024) for zero-shot TTS evaluation; 2) the LibriSpeech-PC test-clean set following F5-TTS (Chen et al., 2024b) for zero-shot TTS evaluation; 3) the L2-arctic dataset (Zhao et al., 2018) following (Melechovsky et al., 2022; Liu et al., 2024a) for accented TTS evaluation.
@@
We train the WaveVAE model and S-DiT on 8 NVIDIA A100 GPUs. The batch sizes, optimizer settings, and learning rate schedules are described in Appendix A.1. It takes 2M steps for the WaveVAE model’s training and 1M steps for S-DiT’s training until convergence. The pre-training of S-DiT requires 800k steps and PeRFlow distillation requires 200k steps.

### v4

# MegaTTS 3: Sparse Alignment Enhanced Latent Diffusion Transformer for Zero-Shot Speech Synthesis

@@ Abstract
While recent zero-shot text-to-speech (TTS) models have significantly improved speech quality and expressiveness, mainstream systems still suffer from issues related to speech-text alignment modeling: 1) models without explicit speech-text alignment modeling exhibit less robustness, especially for hard sentences in practical applications; 2) predefined alignment-based models suffer from naturalness constraints of forced alignments. This paper introduces MegaTTS 3, a TTS system featuring an innovative sparse alignment algorithm that guides the latent diffusion transformer (DiT). Specifically, we provide sparse alignment boundaries to MegaTTS 3 to reduce the difficulty of alignment without limiting the search space, thereby achieving high naturalness. Moreover, we employ a multi-condition classifier-free guidance strategy for accent intensity adjustment and adopt the piecewise rectified flow technique to accelerate the generation process. Experiments demonstrate that MegaTTS 3 achieves state-of-the-art zero-shot TTS speech quality and supports highly flexible control over accent intensity. Notably, our system can generate high-quality one-minute speech with only 8 sampling steps. Audio samples are available at https://sditdemo.github.io/sditdemo/.
MegaTTS 3: Sparse Alignment Enhanced Latent Diffusion Transformer
@@
Ziyue Jiang1,2thanks: Intern at ByteDance., Yi Ren2, Ruiqi Li1,2, Shengpeng Ji1, Boyang Zhang1, Zhenhui Ye1, Chen Zhang2, Bai Jionghao1, Xiaoda Yang1, Jialong Zuo1, Yu Zhang1, Rui Liu3, Xiang Yin2, Zhou Zhao1 1Zhejiang University, 2ByteDance, 3Inner Mongolia University ziyuejiang341@gmail.com, zhaozhou@zju.edu.cn
@@ 1 Introduction
In recent years, neural codec language models (Wang et al., 2023; Zhang et al., 2023; Song et al., 2024; Xin et al., 2024) and large-scale diffusion models  (Shen et al., 2023; Matthew et al., 2023; Lee et al., 2024a; Eskimez et al., 2024; Ju et al., 2024; Yang et al., 2024d, b) have brought considerable advancements to the field of speech synthesis. Unlike traditional text-to-speech (TTS) systems (Shen et al., 2018; Jia et al., 2018; Li et al., 2019; Kim et al., 2020; Ren et al., 2019; Kim et al., 2021, 2022), these models are trained on large-scale, multi-domain speech corpora, which contributes to notable improvements in the naturalness and expressiveness of synthesized audio. Given only seconds of speech prompt, they can synthesize identity-preserving speech in a zero-shot manner.
@@
  • Predefined alignment-based methods have prosodic naturalness constraints of forced alignments. During training, alignment paths (Ren et al., 2020; Kim et al., 2020) are directly introduced into their models (Matthew et al., 2023; Shen et al., 2023; Ju et al., 2024) to reduce the complexity of text-to-speech generation, which achieves higher intelligibility. Nevertheless, they suffer from the following limitations: 1) predefined alignments constrain the model’s search space to produce more natural-sounding speech (Anastassiou et al., 2024; Chen et al., 2024a); 2) the overall naturalness is highly dependent on the performance of duration models.
Intuitively, we can integrate the two aforementioned diffusion-based methods to pursue optimal performance. To be specific, we propose a novel sparse speech-text alignment strategy to enhance the latent diffusion transformer (DiT), termed MegaTTS 3. In our approach, phoneme tokens are sparsely distributed within the corresponding forced alignment regions to provide coarse pronunciation information that is then refined by the latent DiT model. Experimental results demonstrate that MegaTTS 3 achieves nearly state-of-the-art speech intelligibility and speaker similarity on the LibriSpeech test-clean set (Panayotov et al., 2015) with only 8 sampling steps, while also exhibiting high speech naturalness. The main contributions of this work are summarized as follows:
  • We design a sparse alignment enhanced latent diffusion transformer model, which effectively integrates the strengths of the two aforementioned speech-text alignment approaches. Notably, our model also demonstrates greater robustness to duration prediction errors compared to methods with forced alignment.
@@
  • We successfully reduce the inference steps from 25 to 8 with the piecewise rectified flow (PeRFLow) technique, achieving highly efficient zero-shot TTS with minimal quality degradation. We also visualize the attention matrices across various layers of MegaTTS 3 and obtain insightful findings in Appendix F.
@@ 2 Background
Figure 1: (a) The WaveVAE model; (b) Overview of our model. We insert the sparse alignment anchors into the latent vector sequence to provide coarse alignment information. The transformer blocks in MegaTTS 3 will automatically build fine-grained alignment paths.
@@ 3 Method
This section introduces MegaTTS 3. To begin with, we describe the architecture design of MegaTTS 3. Then, we provide detailed explanations of the sparse alignment mechanism, the piecewise rectified flow acceleration technique, and the multi-condition classifier-free guidance strategy.
@@ 3.1 Architecture
We use the standard transformer block from LLAMA (Dubey et al., 2024) as the basic structure for MegaTTS 3 and adopt the Rotary Position Embedding (RoPE) (Su et al., 2024) as the positional embedding. During training, we randomly divide the latent vector sequence into a prompt region $z_{prompt}$ and a masked target region $z_{target}$, with the proportion of $z_{prompt}$ being $\gamma\sim U(0.1,0.9)$. We use $v_{\theta}$ to predict the masked target vector $\hat{z}_{target}$ conditioned on $z_{prompt}$ and the phoneme embedding $p$, denoted as $v_{\theta}(\hat{z}_{target}|z_{prompt},p)$. The loss is calculated using only the masked region $z_{target}$. MegaTTS 3 learns the average pronunciation from $p$ and the specific characteristics such as timbre, accent, and prosody of the corresponding speaker from $z_{prompt}$.
3.2 Sparse Alignment Enhanced Latent Diffusion Transformer (MegaTTS 3)
In this subsection, we describe the sparse alignment strategy as the foundation of MegaTTS 3, followed by the piecewise rectified flow and multi-condition CFG strategies to further enhance MegaTTS 3’s capacity.
@@ 3.2 Sparse Alignment Enhanced Latent Diffusion Transformer (MegaTTS 3)
Let’s first analyze the reasons behind the characteristics of different speech-text alignment modeling methods in depth. Implicitly modeling speech-text alignment is a relatively challenging task, which consequently leads to suboptimal speech intelligibility, particularly in hard cases. On the other hand, employing predefined hard alignment paths constrains the model’s search space to produce more natural-sounding speech. The characteristics of these systems motivate us to design an approach that combines the advantages of both: we first provide a rough alignment to MegaTTS 3 and then use attention mechanisms in Transformer blocks to construct the fine-grained implicit alignment path. The visualizations of the implicit alignment paths are included in Appendix F. In specific, denote the latent speech vector sequence as $z=[z_{1},z_{2},\cdots,z_{n}]$, the phoneme sequence as $p=[p_{1},p_{2},\cdots,p_{m}]$, and the phoneme duration sequence as $d=[d_{1},d_{2},\cdots,d_{m}]$, where $n$, $m$ is the length of the sequence. The length of the speech vector that corresponds to a phoneme $p_{i}$ is the duration $d_{i}$. Given $d=[2,2,3]$, the hard speech-text alignment path can be denoted as $a=[p_{1},p_{1},p_{2},p_{2},p_{3},p_{3},p_{3}]$. To construct the rough alignment $\tilde{a}$, we randomly retain only one anchor for each phoneme: $\tilde{a}=[\underline{M},p_{1},p_{2},\underline{M},\underline{M},\underline{M},P_{3}]$, where $\underline{M}$ represents the mask token. $\tilde{a}$ is downsampled with convolution layers to match the length of the latent sequence $z$. Then, we directly concatenate the downsampled $\tilde{a}$ and $z$ along the channel dimension. The anchors in $\tilde{a}$ provide MegaTTS 3 with approximate positional information for each phoneme, simplifying the learning process of speech-text alignment. At the same time, the rough alignment information does not limit MegaTTS 3’s search space and also enables fine-grained control over each phoneme’s duration.
@@
We adopt Piecewise Rectified Flow (PeRFlow) (Yan et al., 2024) to distill the pretrained MegaTTS 3 model into a more efficient generator. Although our MegaTTS 3 is non-autoregressive in terms of the time dimension, it requires multiple iterations to solve the Flow ODE. The number of iterations (i.e., number of function evaluations, NFE) has a great impact on inference efficiency, especially when the model scales up further. Therefore, we adopt the PeRFlow technique to further reduce NFE by segmenting the flow trajectories into multiple time windows. Applying reflow operations within these shortened time intervals, PeRFlow eliminates the need to simulate the full ODE trajectory for training data preparation, allowing it to be trained in real-time alongside large-scale real data during the training process. Given number of windows $K$, we divide the time $t\in[0,1]$ into $K$ time windows $\{(t_{k-1},t_{k}]\}^{K}_{k=1}$. Then, we randomly sample $k\in\{1,\cdots,K\}$ uniformly. We use the startpoint of the sampled time window $z_{t_{k-1}}=\sqrt{1-\sigma^{2}(t_{k-1})}z_{1}+\sigma(t_{k-1})\epsilon$ to solve the endpoint of the time window $\hat{z}_{t_{k}}=\phi_{\theta}(z_{t_{k-1}},t_{k-1},t_{k})$, where $\epsilon\sim\mathcal{N}(0,I)$ is the random noise, $\sigma(t)$ is the noise schedule, and $\phi_{\theta}$ is the ODE solver of the teacher model. Since $z_{t_{k-1}}$ and $\hat{z}_{t_{k}}$ is available, the student model $\hat{\theta}$ can be trained via the following objectives:
@@
where $v_{\hat{\theta}}$ is the estimated drift force with parameter $\hat{\theta}$ and $t$ is uniformly sampled from $(t_{k-1},t_{k}]$. We provide details of PeRFlow training for MegaTTS 3 in Appendix C.
@@
where $c$ denotes the conditional state, $\varnothing$ denotes the unconditional state, and $\alpha$ is the guidance scale selected based on experimental results. Unlike standard classifier-free guidance, MegaTTS 3’s conditional states $c$ consist of two components: phoneme embeddings $p$ and the speaker prompt $z_{prompt}$. In the experiments, as the text guidance scale increases, we observe that the pronunciation changes according to the following pattern: 1) starting with improper pronunciation; 2) then shifting to pronouncing with the current speaker’s accent; 3) and finally approaching the standard pronunciation of the target language. The detailed experimental setup is described in Appendix K. This allows us to use the text guidance scale $\alpha_{txt}$ to control the accent intensity. At the same time, the speaker guidance scale $\alpha_{spk}$ should be a relatively high value to ensure a high speaker similarity. Therefore, we adopt the multi-condition classifier-free guidance technique to separately control $\alpha_{txt}$ and $\alpha_{spk}$:
@@
MegaTTS 3 0.3B LibriLight 0.71 0.78 1.82% 0.00 3.98 0.188
MegaTTS 3-accelerated 0.3B LibriLight 0.70 0.78 1.86% -0.03 3.96 0.124
@@ 4.1 Experimental setup
We train MegaTTS 3 on the LibriLight (Kahn et al., 2020) dataset, which contains 60k hours of unlabeled speech derived from LibriVox audiobooks. All speech data are sampled at 16KHz. We transcribe the speeches using an internal ASR system and extract the predefined speech-text alignment using the external alignment tool (McAuliffe et al., 2017). We utilize three benchmark datasets: 1) the librispeech (Panayotov et al., 2015) test-clean set following NaturalSpeech 3 (Ju et al., 2024) for zero-shot TTS evaluation; 2) the LibriSpeech-PC test-clean set following F5-TTS (Chen et al., 2024b) for zero-shot TTS evaluation; 3) the L2-arctic dataset (Zhao et al., 2018) following (Melechovsky et al., 2022; Liu et al., 2024a) for accented TTS evaluation.
@@
We train the WaveVAE model and MegaTTS 3 on 8 NVIDIA A100 GPUs. The batch sizes, optimizer settings, and learning rate schedules are described in Appendix A.1. It takes 2M steps for the WaveVAE model’s training and 1M steps for MegaTTS 3’s training until convergence. The pre-training of MegaTTS 3 requires 800k steps and PeRFlow distillation requires 200k steps.
@@
4.2 Results of Zero-Shot Speech Synthesis
Evaluation Baselines.
We compare the zero-shot speech synthesis performance of MegaTTS 3 with 11 strong baselines, including: 1) VALL-E 2 (Chen et al., 2024a); 2) VoiceBox (Matthew et al., 2023); 3) DiTTo-TTS (Lee et al., 2024a); 4) NaturalSpeech 3 (Ju et al., 2024); 5) CosyVoice (Du et al., 2024); 6) MaskGCT (Wang et al., 2024); 7) F5-TTS (Chen et al., 2024b); 8) E2 TTS (Eskimez et al., 2024). Explanation and details of the selected baseline systems are provided in Appendix A.4.
Analysis
As shown in Table 1, we can see that 1) MegaTTS 3 achieves state-of-the-art SIM-O, SMOS, and WER scores, comparable to NaturalSpeech 3 (the counterpart with forced alignment), and significantly surpasses other baselines without explicit alignments. The improved SIM-O and SMOS suggest that the proposed sparse alignment effectively simplifies the text-to-speech mapping challenge like predefined forced duration information, allowing the model to focus more on learning timbre information. And the improved WER indicates that MegaTTS 3 also enjoys strong robustness; 2) MegaTTS 3 significantly surpasses all baselines in terms of CMOS, demonstrating the effectiveness of the proposed sparse alignment strategy; 3) After the PeRFlow acceleration, the student model of MegaTTS 3 shows on par quality with the teacher model and enjoys fast inference speed. We also conduct the experiments on the LibriSpeech-PC test-clean set provided by F5-TTS and the results are shown in Table 2, which also demonstrates that our method achieves state-of-the-art performance in terms of speaker similarity and speech intelligibility. The duration controllability of MegaTTS 3 is verified in Appendix E. In the demo page, we also demonstrate that our method can maintain high naturalness even when the performance of the duration predictor is suboptimal (while MegaTTS 3 with forced alignment fails).
GT - 45.1 0.591 0.783 4.03 +0.09 3.95
CTA-TTS 5.98 41.1 0.602 0.799 3.72 -0.60 3.64
MegaTTS 3 5.69 42.3 0.601 0.790 3.84 +0.00 3.89